ACCURATE POTENTIAL ENERGY CURVES FOR THE GROUND ELECTRONIC STATES OF NeH^{+} AND ArH^{+}

JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia, Canada B3H 4J3; \underline{PHOTOS G. HAJIGEORGIOU}, Department of Life and Health Sciences, University of Nicosia, Nicosia 1700, Cyprus.

All available microwave and infrared spectroscopic line positions for the ground electronic states of the molecular cations NeH^{+} and ArH^{+} were employed in a direct potential fitting procedure to determine compact analytical potential curves and radial functions describing breakdown of the Born-Oppenheimer approximation. For NeH^{+}, 17 adjustable parameters were required to represent a total of 183 line positions for 4 isotopologues, whereas for ArH^{+}, 23 adjustable parameters were required to represent 440 line positions for 6 isotopologues. The MLR3 potential energy functional form was employed, taking full account of the proper 1/r^{4} limiting long-range dependence of the ion-atom dispersion energy interactions. Accurate vibrational energies, rotational constants and centrifugal distortion constants have been calculated for both diatomic cations.